. Superphosphate, or sometimes called triple phosphate, is a commonly used water-soluble fertilizer. It is produced by the reaction: πΆπ‘Ž3 (𝑃𝑂4 )2 + 𝐻2𝑆𝑂4 β†’ πΆπ‘Ž(𝐻2𝑃𝑂4 )2 + πΆπ‘Žπ‘†π‘‚4 If we mix 200.0 g of πΆπ‘Ž3 (𝑃𝑂4 )2 with 133.5 g of 𝐻2𝑆𝑂4, how much πΆπ‘Ž(𝐻2𝑃𝑂4 )2 (in grams) can be produced?


1
Expert's answer
2020-09-05T14:41:01-0400

The Chemical equation is,

"Ca_3(PO_4)_2\ +H_2SO_4\ \to Ca(H_2PO_4)_2\ +CaSO_4\ where,\newline Ca_3(PO_4)_2=200g\ and,\newline H_2SO_4=133.5g"


A balanced chemical equation of the reaction is

"Ca_3(PO_4)_2\ +2H_2SO_4\ \to\ Ca(H_2PO_4)_2\ +\ 2CaSO_4"


Calculate the number of moles of each reagent.

"moles=\dfrac{mass}{molar\ mass}"


"Ca_3(PO_4)_2\ mols\ =\dfrac{200}{310.18}\ =0.6448\ mol"


"H_2SO_4\ mols=\dfrac{133.5}{98.079}\ = 1.361\ mol"


Express moles of each reagent in terms moles of the product

"Ca(H_2PO_4)_2\ moles=0.6448*\dfrac{1 molCa(H_2PO_4)_2}{1\ molCa_3(PO_4)_2}"

"=0.6448mol"


"Ca(H_2PO_4)_2\ moles=1.361*\dfrac{1\ molCa(H_2PO_4)_2}{2moles H_2SO_4}"

"=0.6805mol"


Since this is limiting reagent problem, the reagent that produces a smaller number of moles of the product is used in calculating the mass of the product.In this case "Ca_3(PO_4)_2" has the smaller number of moles .


"Mass=no.moles\ of\ Ca_3(PO_4)_2*\ molar \ massCa(H_2PO_4)_2"

"=0.6448*234.1\newline =150.9g"